# Identifying Spurious Correlations using Counterfactual Alignment

### v1

@@ Abstract
Models driven by spurious correlations often yield poor generalization performance. We propose the counterfactual alignment method to detect and explore spurious correlations of black box classifiers. Counterfactual images generated with respect to one classifier can be input into other classifiers to see if they also induce changes in the outputs of these classifiers. The relationship between these responses can be quantified and used to identify specific instances where a spurious correlation exists as well as compute aggregate statistics over a dataset. Our work demonstrates the ability to detect spurious correlations in face attribute classifiers. This is validated by observing intuitive trends in a face attribute classifier as well as fabricating spurious correlations and detecting their presence, both visually and quantitatively. Further, utilizing the CF alignment method, we demonstrate that we can rectify spurious correlations identified in classifiers.
00footnotetext: 1Work done prior to joining Amazon
@@ 1 Introduction
Challenges related to neural network generalization and fairness often arise due to covariate shift [19] and shortcut learning [10]. Shortcut learning can lead to models making decisions based on factors not aligned with expectations of the model creator. These powerful models (vision models in this work) leverage a wide array of features and relationships, which may inadvertently incorporate undesired, spurious correlations. These spurious relationships may stem from sample bias (e.g., predicting cows when cows are observed on grass but not on a beach) or may be inherent to the class definition (e.g., predicting cows when an animal has four legs) [6].
Our primary goal is to ensure that models make predictions without relying on spurious correlates, a concept coined as “right for the right reasons” [25]. This means that the features employed for prediction align with expectations of the model creator, rather than features which are spuriously correlated with the target label. Models driven by such spurious correlations often yield poor generalization performance [6, 10]. Consequently, it is crucial to develop methods for uncovering and correcting these unwanted correlations.
Our objective is to understand black-box classifiers, such as pre-trained models, as this is a common use case encountered by practitioners. In this analysis, we utilize counterfactual (CF) images, synthetic images simulating a change in the class label of an image [35]. These synthetic images have features modified such that the prediction of the classifier changes. We can then view the synthetic images to understand the reasons that a prediction was made.
Specifically, we are interested in CF images that are directly generated using the gradients of a classifier [8, 12, 3]. Generating CF images with respect to a classifier is rooted in similar logic to that of crafting adversarial examples. However, the key distinction lies in the constraint that CF images remain within the data manifold of plausible images. The latent space of an autoencoder provides such data manifold. This approach enables us to study the specific features used by a given classifier on a particular input. By keeping the classifier and autoencoder independent, different classifiers can be analysed using a singular fixed autoencoder as a reference point.
@@
The approach we take is to generate CF images for the positive predictions of a classifier over an entire dataset and study the resulting CF images using a collection of different classifiers, which we refer to as downstream classifiers, to observe what outputs are impacted. If the CF images generated for one classifier also change the predictions of other classifiers, we can conclude there is shared feature usage. This shared feature usage can prompt us to investigate unexpected relationships. The relative change metric (Eq. 3) can be used to identify spurious correlations which may occur for a specific example or to identify trends over an entire dataset. We study face attribute classification, recognizing its advantage in allowing visual verification by readers.
@@
  • We propose CF alignment to reason about the feature relationships between classifiers. This approach allows for both aggregate quantification and targeted querying of models to identify instances where incorrect reasoning leads to a prediction.
  • Our work demonstrates the ability to detect spurious correlations in face attribute classifiers. This is validated by observing intuitive trends in a face attribute classifier as well as inducing spurious correlations and then detecting their presence, both visually and quantitatively.
  • Utilizing the CF alignment method, we demonstrate that we can rectify spurious correlations identified in classifiers. Experimental results show a bias reduction in 10 out of 12 classifiers.
@@ 2 Related Work
Counterfactual (CF) generation can be done in a variety of ways [35]. Classifier specific approaches generally perturb a latent space representation guided by a classifier. Some methods use the gradient of a classifier to guide movement in the latent space by computing the gradient directly [8] or defining a loss that is optimized [12, 3]. Another classifier based approach is to train a model that predicts, using training examples, where to walk in the latent space to change the classifier’s prediction [28]. This has the benefit of working with classifiers that are not differentiable. Another approach is conditional generation which doesn’t use a classifier and instead generates images based on a conditioning variable. This conditioning can be a label [18, 5, 27, 28, 11, 4, 31] or text [7] or by manually adjusting the latent representation [29]. Conditional generation learns a representation from the data and does not capture the exact features utilized by a classifier, as the gradient based methods do.
@@
Several methods exist for studying neural network based classifiers to understand the features they use. Early approaches include gradient based attribute maps (aka saliency heatmaps) that could be overlaid on the image [30]. This was extended to capture more caveats of neural network reasoning[33, 32] but these methods still only generate heatmap explanations. Other work focuses on leveraging occlusions to identify discrete image regions [23]. Other work inspects individual neural network neurons to identify where class information was propagating using linear probing [1] and manual identification [21, 20]. Additional methods focus on concepts represented by layer activations and identifying images that share these representations [14]. Another approach [2] generates synthetic images using human annotators such that the images have known feature similarity (e.g. a face with all features held constant except for skin color). These images are then used to identify spurious correlations between attributes.
3 Alignment Methodology
This work proposes the CF alignment approach as described in Fig. 1. For this approach, we generate counterfactuals using Latent Shift [8]. This approach requires an encoder/decoder model $D(E(x))$ where $E$ is the encoder and $D$ is the decoder, and a classifier $f$ which predicts a target $y$ as follows: $y=f(x)$. The autoencoder and the classifier are trained independently without specific requirements except for being differentiable and operating on the same data domain. To compute a counterfactual, an input image $x$ is encoded using $E(x)$, to produce a latent representation $z$. Perturbations of the latent space are computed for a base classifier $f_{b}$ in Eq. 1 which is then used to produce $\lambda$-shifted samples as shown in Eq. 2. $\lambda$ is determined using an iterative search where the value is decreased in steps until the classifier’s prediction is reduced by 0.6 or begins to increase. $z_{\lambda}=z-\lambda\frac{\partial f_{b}(D(z))}{\partial z}$ (1) $x^{\prime}_{\lambda}=D(z_{\lambda})$ (2)
These counterfactual samples can then be input into downstream classifiers that predict various attributes (e.g. $f_{1}$, $f_{2}$) and observe how their predictions change. Relative change (Eq. 3) allows us to quantify the relationship between a base and downstream classifier, similar to correlation, but accounts for the magnitude of the change. In our experiments, we found correlation would produce false positives when there was only a slight change in the prediction of $f_{1}$ compared to the base classifier $f_{b}$.
$\frac{f_{1}(D(z_{\lambda}))-f_{1}(D(z_{0}))}{f_{b}(D(z_{\lambda}))-f_{b}(D(z_{0}))}$ (3)
@@ 4 Experiments
The experiments in this work are performed on the CelebA dataset [17] that contains over 200k celebrity images with 40 facial attribute labels per image. The resolution of the images is 1024$\times$1024.
The pre-trained face classifiers used in this work are sourced from [34]. They were trained on the CelebA dataset [17] to predict 40 different facial attributes on images of dimensions 224$\times$224. We chose this model because it has good performance, is publicly available, and was implemented in PyTorch.
We leverage the VQ-GAN autoencoder from [9] trained on the FacesHQ dataset, which combines the CelebA dataset [17] and the Flickr-Faces-HQ (FFHQ) dataset [13]. The resolution of this model is 256x256.
The Captum library [15] is utilized for baseline attribution methods. PyTorch [22] is used for efficient tensor computation and automatic differentiation. The source code and model weights for all experiments will be released publicly online [^1] https://github.com/ieee8023/latentshift.
Figure 2: Relationships between face attribute classifiers as measured by CF-induced relative changes (left), classifier predictions (middle) and training data labels (right). In (a), base classifiers are along the rows and downstream classifiers are along the columns. Comparing (a) to (b) and (c), shows that many relationships reflected in the CF outputs are preserved from correlations in the training data. The non-symmetric relationship between male and wearing_lipstick, highlighted in green, reflects expected interactions between the two face attribute classifiers. The relationship between male and big_nose, highlighted in red, is strong in both the classifier predictions and ground truth labels but low in CF alignment, indicating that although correlated, these features aren’t exploited by the classifier.
(a) CF alignment for male showing alignment with classifiers related to hair and inverse alignment with classifiers related to makeup. The CF image visually appears to add makeup and remove hair.
(b) CF alignment for attractive showing alignment with classifiers related to makeup, face shape, and age and an inverse alignment with nose size. The CF image visually appears to remove makeup and add bags under the eyes.
(c) CF alignment for heavy_makeup showing alignment with classifiers related to makeup and face shape and inverse alignment with classifiers for nose size and bags under the eyes. The CF image visually shows makeup being removed.
Figure 3: Examples of alignment for specific samples showing the visual counterfactual changes as well as other classifiers with aligned and inverse aligned relationships.
@@ 4.1 Aggregate statistics over a dataset
Viewing aggregate CF alignment statistics over a dataset can be useful when investigating a model for bias or spurious correlations. The average relative change between pairs of classifiers for 400 samples per class is shown in Fig. 2a. Figures 2b and 2c show correlations between classifier predictions and the ground truth training labels. Only an interesting subset of classifiers are shown in this figure for clarity, with the full CF alignment matrix in Appendix Fig. 9.
We draw the readers attention to the example relationship between male and big_nose in red, which is high in both classifier prediction and ground truth labels but very low using CF alignment. This indicates that while the attributes (labels) are correlated, the male classifier does not largely use big_nose as a feature to predict male - an expected insight that is verified. A similar insight is found for the class of mouth_slightly_open, which has a relatively strong relationship with smiling.
We also observe many non-symmetric relationships such as male and wearing_lipstick in green, as well as young and attractive. These are also intuitive relationships that we would expect the classifiers to learn. The male classifier likely uses lipstick as a feature to reduce the likelihood that the image contains a male subject (high inverse correlation in Fig. 2c), while the wearing_lipstick classifier would be more focused on the color of the lips. Overall, these results help qualitatively depict that counterfactual reasoning can provide model introspection in a manner that is consistent with human reasoning.
@@ 4.2 Studying specific examples
Using the aggregate analysis from §4.1 as a guide, classifiers can be selected to investigate undesired relationships. We will first select a sample which contains an inverse relationship between a base classifier for male and downstream classifier for heavy_makeup in Fig. 2(a). Visually, we can see makeup and lipstick in the CF image and the predictions increase as expected in the classifiers for heavy_makeup, wearing_lipstick, and no_beard. Similarly, there is a decrease in the classifier for goatee. These appear reasonable and provide insight about the features used for prediction. Exploring the more subjective base classifier label attractive in Fig. 2(b) we observe big_nose and bags_under_eyes increase, which can depict human preferences and biases in determining an attractive face.
@@
4.3 Comparison with saliency maps
Saliency maps for feature attribution could potentially be used to perform a similar analysis by looking at relationships between their generated heatmaps. Here, we present an example that demonstrates their limitation. The attribution methods input gradients, integrated gradients [33], and guided backprop [32] are used to explain the classifier predictions of pointy_nose and big_nose. In Fig. 4 the salient areas for pointy_nose and big_nose are the same and cannot be disambiguated. Comparing these heatmaps would yield a high similarity but we observe in Fig. 2 that these two classifiers have an inverse relationship. Using CF alignment, we can gain a deeper understanding of what features the model is using and better reason about why the decision was made.

### v2

@@ Abstract
Models driven by spurious correlations often yield poor generalization performance. We propose the counterfactual (CF) alignment method to detect and quantify spurious correlations of black box classifiers. Our methodology is based on counterfactual images generated with respect to one classifier being input into other classifiers to see if they also induce changes in the outputs of these classifiers. The relationship between these responses can be quantified and used to identify specific instances where a spurious correlation existsThThis is validated by observing intuitive trends in a face-attribute face-attribute and waterbird classifiers, as well as by fabricating spurious correlations and detecting their presence, both visually and quantitatively. Furthermore, utilizing the CF alignment method, we demonstrate that we can evaluate robust optimization methods (GroupDRO, JTT, and FLAC) by detecting a reduction in spurious correlations.
00footnotetext: AWork not related to position at Amazon.
@@ 1 Introduction
Challenges related to neural network generalization and fairness often arise due to covariate shift (Moreno-Torres et al., 2012) and shortcut learning (Ross et al., 2017; Geirhos et al., 2020). Shortcut learning can lead to models making decisions based on factors not aligned with expectations of the model creator. These powerful models (vision models in this work) leverage a wide array of features and relationships, which may inadvertently incorporate unwanted spurious correlations. These spurious relationships may stem from sample bias (e.g., predicting cows when cows are observed on grass but not on a beach) or may be inherent to the class definition (e.g., predicting cows when an animal has four legs) (Beery et al., 2018).
In this work our objective is to understand black-box classifiers, without access to their training data, as this is a common use case encountered by practitioners. In this analysis, we utilize counterfactual (CF) images which are synthetic images simulating a change in the class label of an image (Pearl, 2009). These synthetic images have features modified such that the prediction of the classifier changes. We can then view the synthetic images to understand the reasons that a prediction was made.
Specifically, we are interested in CF images that are directly generated using the gradients of a classifier (Cohen et al., 2021; Joshi et al., 2018; Balasubramanian et al., 2020). Generating CF images with respect to a classifier is rooted in similar logic to that of crafting adversarial examples. However, the key distinction lies in the constraint that CF images remain within the data manifold of plausible images. The latent space of an autoencoder provides such data manifold. This approach enables us to study the specific features used by a given classifier on a particular input. By keeping the classifier and autoencoder independent, different classifiers can be analysed using a singular fixed autoencoder as a reference point.
@@
The approach we take is to generate CF images for the positive predictions of a classifier over an entire dataset and study the resulting CF images using a collection of different classifiers, which we refer to as downstream classifiers, to observe what outputs are impacted. If the CF images generated for one classifier also change the predictions of other classifiers, we can conclude that there is shared feature usage. This shared feature usage can prompt us to investigate unexpected relationships. The relative change metric (Eq. 3) can be used to quantify the impact of counterfactual perturbations and identify spurious correlations that may occur for a specific example or to identify trends across an entire dataset. We study face-attribute classification, recognizing its advantage in allowing visual verification by readers.
@@
  • We propose CF alignment to reason about the feature relationships between classifiers, quantified using the relative change metric. This approach allows for both aggregate quantification and targeted querying of models to locate specific examples where the spurious correlations are used to make predictions.
  • Our work demonstrates the ability to detect spurious correlations on existing face-attribute and waterbird classifiers. This is validated by observing intuitive trends in a face-attribute classifier as well as inducing spurious correlations and then detecting their presence, both visually and quantitatively.
  • We demonstrate that CF Alignment can evaluate robust optimization methods (GroupDRO, JTT, and FLAC) by detecting a reduction in spurious correlations. We also observe improved generalization performance when a spurious correlation is reduced.
@@ 2 Related Work
Counterfactual (CF) generation can be done in a variety of ways (Verma et al., 2020). Classifier specific approaches generally perturb a latent space representation guided by a classifier. Some methods use the gradient of a classifier to guide movement in the latent space by computing the gradient directly (Cohen et al., 2021) or defining a loss that is optimized (Joshi et al., 2018; Balasubramanian et al., 2020). Another classifier based approach is to train a model that predicts, using training examples, where to walk in the latent space to change the classifier’s prediction (Schutte et al., 2020). This has the benefit of working with classifiers that are not differentiable. Another approach is conditional generation which doesn’t use a classifier and instead generates images based on a conditioning variable. This conditioning can be a label (Mirza & Osindero, 2014; Baumgartner et al., 2018; Samangouei et al., 2018; Schutte et al., 2020; Hasenstab et al., 2023; Barredo-Arrieta & Del Ser, 2020; Singla et al., 2023) or text (Chambon et al., 2022), or can be provided by manually adjusting the latent representation (Seah et al., 2019). Conditional generation learns a representation from the data and does not capture the exact features utilized by a classifier, as the gradient based methods do.
@@
Several methods exist for studying neural network based classifiers to understand the features they use. Early approaches include gradient based attribute maps (aka saliency heatmaps) that could be overlaid on the image (Simonyan et al., 2014). This was extended to capture more caveats of neural network reasoning (Sundararajan et al., 2017; Springenberg et al., 2015) but these methods still only generate heatmap explanations. Other work focuses on leveraging occlusions to identify discrete image regions (Ribeiro et al., 2016) that are relevant to a prediction. This is useful when predictions are based on discrete features that are smaller than the occlusion size. Other work inspects individual neural network neurons to identify where class information is propagating using linear probing (Alain & Bengio, 2016) and manual identification (Olah et al., 2020; 2017). This is useful for understanding the internal representations of neural networks but does not explain individual predictions.
Along these lines, work by Kim et al. (2018) focuses on concepts represented by layer activations forming so called “Concept Activation Vectors”. These vectors allow us to identify inputs which result in a similar internal state and therefore have a similar reason for being predicted. A limitation of this approach is that it is difficult to compare vectors between classifiers because they have different weights.
An approach by Balakrishnan et al. (2021) generates synthetic images using human annotators such that the images have known feature similarity (e.g. a face with all features held constant except for skin color). These images can then be used to identify spurious correlations between attributes. To contrast this approach to ours, this work is creating CFs manually using humans and then looking for unexpected changes in classifier predictions, similar to our approach. A limitation of this method is the required manual effort and the reliance on human knowledge of concepts. The automatic generation of CFs in our work overcomes this issue and allows us to identify features that are unknown initially to humans. And the relative change metric makes it easy to identify potential spurious correlations.
3 CF Alignment Methodology
In this work, we introduce the CF alignment approach, outlined in Figure 1. The methodology involves the generation of counterfactual samples using a technique known as Latent Shift (Cohen et al., 2021). The implementation of Latent Shift requires the integration of an encoder/decoder model, denoted as $D(E(x))$, where $E$ represents the encoder and $D$ is the decoder. Additionally, a classifier $f$ is incorporated, responsible for predicting the target variable $y$, expressed as $y=f(x)$. It is important to note that both the autoencoder and the classifier are trained independently, with the only specified requirements being differentiability and operation on the same data domain.
To compute a counterfactual, the process begins with encoding an input image $x$ using the encoder $E(x)$, resulting in a latent representation $z$. The next step involves perturbing the latent space to generate counterfactual samples. This perturbation is performed using a base classifier $f_{b}$, as illustrated in Equation 1. The resulting perturbed latent representation is denoted as $z_{\lambda}$. Subsequently, the decoder $D$ is employed to reconstruct the image, resulting in a counterfactual image $x^{\prime}_{\lambda}$, as depicted in Equation 2.
The perturbation of the latent space, represented by $z_{\lambda}$, is computed by subtracting $\lambda$ times the gradient of the base classifier $f_{b}$ with respect to the latent representation $z$. This operation is conducted to induce changes in the latent space that influence the predictions of the classifier. The parameter $\lambda$ is determined through an iterative search process, where its value is systematically adjusted in steps. The objective is to find a suitable $\lambda$ such that the classifier’s prediction is either reduced by 0.6 or starts to increase. 0.6 is chosen as a difference that should cross the decision boundary. Just crossing the 0.5 mark is not always large enough to generate a reasonable counterfactual.
$z_{\lambda}=z_{0}-\lambda\frac{\partial f_{b}(D(z_{0}))}{\partial z}$ (1)
$x^{\prime}_{\lambda}=D(z_{\lambda})$ (2)
Equation 1 expresses the computation of $z_{\lambda}$, and Equation 2 outlines the generation of the counterfactual image $x^{\prime}_{\lambda}$ by decoding the perturbed latent representation.
This process allows us to systematically explore and manipulate the latent space to generate counterfactual samples that reveal insights into the decision-making process of the classifier and its sensitivity to changes in the input data.
3.1 Relative Change
Having generated counterfactual samples using the Latent Shift approach, we proceed to algorithmically assess the impact of these samples on various downstream classifiers responsible for predicting the probability of different attributes, denoted as $f_{1},f_{2},$ and so on. These classifiers are distinct from the base classifier, denoted as $f_{b}$, which was utilized in the counterfactual generation process.
To quantify the relationship between the base classifier and each downstream classifier, we employ the “relative change metric”, as defined in Equation 3. This metric is akin to correlation but takes into account not only the direction of change in predictions but also the magnitude of that change. The formula for relative change is expressed as follows:
$R(f_{1},f_{b},z_{0})=\frac{f_{1}(D(z_{\lambda}))-f_{1}(D(z_{0}))}{f_{b}(D(z_{\lambda}))-f_{b}(D(z_{0}))}$ (3)
Here, $D(z_{\lambda})$ and $D(z_{0})$ represent the reconstructions of the latent representations $z_{\lambda}$ and $z_{0}$ (perturbed and original, respectively) by the decoder $D$. The numerator captures the change in prediction made by the downstream classifier $f_{1}$ in response to the counterfactual perturbation, while the denominator corresponds to the change in prediction made by the base classifier $f_{b}$ due to the same perturbation.
In our experiments, we opted for relative change over traditional correlation measures. This decision was motivated by the observation that correlation metrics occasionally yielded false positives when only a slight change in the prediction of $f_{1}$ occurred compared to the base classifier $f_{b}$. Relative change provides a more nuanced understanding by considering not just the direction but also the magnitude of the change, offering a more robust assessment of the impact of counterfactual perturbations on downstream classifiers.
The relative change metric is a key contribution that allows us to make sense of the CF alignment results at scale over large datasets.
@@ 4 Experiments
The experiments in this work are performed on the CelebA HQ dataset (Karras et al., 2018) that contains over 200k celebrity images with 40 facial attribute labels per image. The resolution of the images is 1024$\times$1024. Experiments are also performed on the lower resolution (178 x 178) CelebA dataset (Liu et al., 2015).
The pre-trained face classifiers used in this work are sourced from (Vandenhende et al., 2020). They were trained on the CelebA dataset (Liu et al., 2015) to predict 40 different facial attributes on images of dimensions 224$\times$224. We chose this model because it has good performance, is publicly available, and was implemented in PyTorch.
We leverage the VQ-GAN autoencoder from (Esser et al., 2021) trained on the FacesHQ dataset, which combines the CelebA HQ dataset (Karras et al., 2018) and the Flickr-Faces-HQ (FFHQ) dataset (Karras et al., 2021). The resolution of this model is 256x256. In order to make this work for existing classifiers, each classifier normalizes the input image dynamically to match its training domain.
The Captum library (Kokhlikyan et al., 2020) is utilized for baseline attribution methods. PyTorch (Paszke et al., 2019) is used for efficient tensor computation and automatic differentiation. The source code and model weights for all experiments will be released publicly online [^1] https://github.com/ieee8023/latentshift. The CF alignment algorithm requires a few seconds to run for each image on a NVIDIA V100 16GB GPU. The CF generation step is the most expensive and is variable due to the automated search for the optimal lambda. The resulting CF images are then processed by each classifier which scales by the number of classifiers studied.
Figure 2: Relationships between face-attribute classifiers as measured by CF-alignment relative changes (left), classifier predictions (middle) and training data labels (right). In (a), base classifiers are along the rows and downstream classifiers are along the columns. Comparing (a) to (b) and (c), shows that many relationships reflected in the CF outputs are preserved from correlations in the training data. We draw the readers attention to some unique differences. The relationship between male and big_nose, highlighted in red, is strong in both the classifier predictions and ground truth labels but low in CF alignment, indicating that although correlated, these features are not exploited by the classifier. In contrast to this, the relationship between pointy_nose and smiling, highlighted in green, is weak in both the classifier predictions and ground truth labels but high in CF alignment, indicating that this relationship was introduced by the classifier.
@@ 4.1 Aggregate statistics over a dataset
Viewing aggregate CF alignment statistics over a dataset can be useful when investigating a model for bias or spurious correlations. The average relative change between pairs of classifiers (N=400 per class) is shown in Figure 2a. Figures 2b and 2c show correlations between classifier predictions and the ground truth training labels. A subset of classifiers studied are shown in this figure for clarity, with the full CF alignment matrix in Appendix Fig. 7. These classifiers were chosen because they have intuitive and unintuitive relationships with large relative change that help to illustrate the contribution of this method.
The caption of Figure 2 details observations where correlation in the training data and classifier outputs are not the same as the relationships uncovered using CF alignment. These indicate that even if correlated attributes are presented to the classifier it does not cause feature usage to be correlated. These observations also imply that even if uncorrelated attributes are presented to the classifier it may still construct spurious feature relationships.
@@ 4.2 Studying specific examples
Figure 3: CF alignment examples for pointy_nose with the highest aligned and inverse aligned classifiers. We observe an inverse alignment with big_nose and potential spurious relationships with eyebrows, eyes, hair, and smiling. The relative change is shown next to each classifier name.
Using the aggregate analysis from §4.1 as a guide, classifiers can be selected to investigate undesired relationships on specific images. Inspecting the pointy_nose classifier in Figure 3 using two images with positive predictions for pointy_nose we observe similar and contrasting relationships between classifiers. A common theme is the inverse alignment with big_nose which is intuitive as it is the opposite of a pointy nose. rosy_cheeks is a common aligned classifier which does not appear to have an obvious reason and is likely a spurious correlation. An alignment with smiling is only observed in one of the examples.
A takeaway from these examples is the uniqueness of a spurious correlations to specific images with specific features, thus demonstrating the need for the method we present to mine for examples where these spurious correlations can be observed.
4.3 Comparison with saliency maps
@@
Saliency maps for feature attribution could potentially be used to perform a similar analysis by looking at relationships between their generated heatmaps. Here, we present an example that demonstrates their limitation. The attribution methods input gradients, integrated gradients Sundararajan et al. (2017), and guided backprop Springenberg et al. (2015) are used to explain the classifier predictions of pointy_nose and big_nose.
In Fig. 4 the salient areas for pointy_nose and big_nose are the same and cannot be disambiguated. The saliency maps only present information as pixel importance which overlaps because both classifiers use features on the nose. Due to these very similar heatmaps, comparing them would not allow us to conclude that the classifiers are using different features. We observe in Fig. 2 that these two classifiers have an inverse relationship. Using CF alignment, we can gain a deeper understanding of what features the model is using, and we can better reason about why the decision was made.